Suppose you pay 20 dollars to enter into a 500 dollar prize. If you have a 5 percent chance of winning, the expected value of your raffle ticket is?
Answer #1

Expected value of raffle ticket = 0.05(500-20) - 0.95(20) = 24-19= 5

Expected value of raffle ticket is 5